# Geometry-Guided Ray Augmentation for Neural Surface Reconstruction with Sparse Views

### v1

@@ 4.2 Visibility Check
4.3 Radiance Map
After the visibility check, rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$ are ensured to be observed from at least one camera viewpoint. The next question is what’s the proper radiance value of rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$? In previous ray augmentation methods [39], a surface ray has only one known radiance value, since the source view visibility check is not performed, thus they simply assign a color from one input view, which neglects the view-dependent effect and leads to inconsistent pseudo label. However, as shown in the previous discussion, a point in the scene surface could be seen from multiple views and we actually have more information regarding the radiance of surface point $\mathbf{v}$. Thus, a main contribution of our method illustrated as follows is that we utilize all reasonably credible information to reconstruct the radiance distribution of surface point $\mathbf{v}$ to infer the radiance of novel rays. Actually, not only did we fit the radiance distribution but also we achieved extrapolation using spherical harmonics expansion fitting.
@@ 4.3 Radiance Map
$\mathbf{c_{v}}(\mathbf{d})=\sum_{\ell=0}^{\ell_{\mathrm{max}}}\sum_{m=-\ell}^{\ell}k_{\ell}^{m}Y_{\ell}^{m}(\mathbf{d}),$ (4)
@@
$\textbf{S}_{\mathbf{v}}=\{(\mathbf{c}_{j},\mathbf{d}_{j})\}=\{(\textbf{c}_{\textbf{v},j},\mathrm{Norm}(\textbf{o}_{j}-\textbf{v}))\}_{j=1}^{N_{\mathrm{vis}}},$ (5)
@@
$\mathop{\arg\min}\limits_{k_{\ell}^{m}}\>\sum_{j=1}^{N_{vis}}|\mathbf{c_{v}}(\mathbf{d}_{j})-\mathbf{c}_{j}|^{2},$ (6)
@@
Figure 3: Visualization of radiance fitting. We project the radiance value of points distributed on the upper hemisphere into this circle to get a visualization of the fitted radiance distribution. The first row of images depicts the SH fitting results, while the second row showcases the outcomes of interpolation. It reveals that the SH fitting provides more informative radiance prediction results on the top line, while preserving the smoothness of distribution.
@@
Figure 4: Qualitative results of novel view synthesis on DTU dataset [10] and Blender dataset [19]. We train the model using 30% images for real scenes in DTU dataset, and 6% for synthetic scenes in Blender dataset.
5 Experiments
Figure 5: Qualitative comparison of the reconstructed depth and foreground mask. The upper row is the result without novel radiance prediction. The lower row is the result using our method.
@@ 5 Experiments
Interpolation 0.0571 0.0633 0.0101 0.1540
SH Fitting 0.0826 0.0770 0.0897 0.1056
Table 1: MSE in the novel views between interpolation and SH fitting with different ratios between known rays and unknown rays.
Chair
Ficus
Lego
NeuS
MVSNeRF
SparseNeuS
Ours
Figure 6: Qualitative results of mesh reconstruction on Blender dataset [19].
Ratio (Used/All) $\alpha=0.2$ $\alpha=0.3$ $\alpha=0.4$
@@
MVSNeRF 24.65 7.53 24.77 5.67 25.85 5.23
Depth-supervised NeRF 23.41 5.41 22.07 6.41 23.97 6.64
SparseNeuS-ft 24.90 3.70 25.35 3.21 26.82 3.04
NeuS 22.85 5.45 23.09 4.21 23.41 3.41
RayAug (Ours) 26.81 3.58 27.08 3.17 27.59 3.12
Table 2: Quantitative comparison against the other state-of-the-art methods on DTU dataset [10]. We evaluate both novel view synthesis (PSNR) and reconstructed mesh quality (Chamfer Distance) on 20 to 40 percent of the training images. Each training image is randomly chosen.
5.1 Novel Radiance Predicting from Spherical Harmonics
$\mathbf{\hat{c}}=w\mathbf{c}_{0}+(1-w)\mathbf{c}_{1},$ (7)
$w=\frac{d(\mathbf{\hat{v}},\mathbf{v}_{1})}{d(\mathbf{v}_{1},\mathbf{\hat{v}})+d(\mathbf{\hat{v}},\mathbf{v}_{2})},$ (8)
$d(\mathbf{v}_{1},\mathbf{v}_{2})=\arctan\frac{\left|\mathbf{v}_{1}\times\mathbf{v}_{2}\right|}{\mathbf{v}_{1}\cdot\mathbf{v}_{2}},$ (9)
We generate a set of around $10,000$ points with all of their visible colors from the DTU dataset for evaluating the two methods for radiance prediction. The radiance from different views of each point is divided into two training (known) and testing (unknown) views with ratios from $8:2$ to $5:5$, in order to evaluate the performance under the different number of novel views. We report the average normalized Mean Square Error (MSE) in the novel views in Table 1. The average MSE for SH fitting is less than that of interpolation, especially when the number of novel views exceeds the number of known views. A visualization case is shown in Figure 3.
Methods Images Chair Drums Ficus Hot Dog Lego Materials Mic Ship Average
NeuS Ours 6 9
21.41 13.70 20.83 19.74 18.36 14.98 21.72 16.47 20.28
24.59 16.91 23.57 27.73 24.17 18.70 25.97 19.92
22.28 15.69 21.71 22.89 19.96 15.95 23.62 18.62 21.59
27.59 19.37 24.09 26.61 23.10 18.36 27.48 21.77
Table 3: Quantitative comparison on Blender dataset [19]. We compare our method with NeuS [32], and reconstruct 8 scenes using 6 or 9 images. The table shows the PSNR on each reconstructed scene.
5.2 Implementation Details
5.2.1 Dataset
We evaluate our method on the DTU dataset [10] and reconstruct 6 scenes in the dataset. These scenes contain various materials, textures, and complex geometry shapes such as holes, multiple objects, and non-convex parts of objects. Originally, each scene contains 49 or 64 images, and the resolution for training and testing is 1600 $\times$ 1200. We also reconstruct 8 scenes from the Blender dataset [19] using our method. Each scene in the Blender dataset contains 100 training images, and scenes are generated from a physical-based rendering engine and include non-Lambertian effects such as reflections. We leverage the mask provided for training and testing, following the same setting used in [16, 32].
5.2.2 Implementation Details
@@
In general, we implement our ray augmentation strategy based on NeuS [32]. Previous work [20] has shown that the input encoding scheme is important for reconstructing an implicit representation. Thus, we compare and choose the frequency encoding [31] for the position input, and sphere harmonics function encoding [6] for the direction. We find that by replacing the frequency encoding with SH function encoding for direct input, the rendered color can converge faster.
@@
5.2.3 Number of Images
The quality of reconstruction with respect to the number of images given is the main concern of this paper, especially when given a few images. For fair comparison across scenes, we do not directly extract a fixed number of images for each scene like previous paper but leverage the ratio $\alpha$ equal to the number of images used and the number of all images provided in the dataset. To evaluate the performance of view synthesis, we randomly choose 70% as the training set and the rest as the testing test. Among the training images, we further choose $\alpha=\{20\%,30\%,40\%\}$ for both image rendering and surface reconstruction.
5.3 Experiment Results
5.3.1 Quantitative
We conduct quantitative comparisons on image quality and mesh quality with NeuS [32], Depth-supervised NeRF [5], SparseNeuS [16] and MVSNeRF [2] on the DTU dataset. Additionally, we compare our method with NeuS for novel view synthesis task on the Blender dataset. For the comparison, we exploit the official public code of the aforementioned methods and employ Peak Signal-to-Noise Ratio (PSNR) and Chamfer Distance (CD) for measuring image quality and mesh quality, respectively.
In Table 2, our method demonstrates significant improvements in the quality of rendered images for all $\alpha$ compared to NeuS [32]. We also train and fine-tune SparseNeuS [16] following the settings in their paper, and the Chamfer Distance of the reconstructed mesh is better decent since they leverage cross-scene prior knowledge for training, and their method achieves a smaller distance compared to our method when using 40% of the images for training. However, in all other cases, our method reconstructs the most precise mesh as evaluated out of other tested methods. It is worth highlighting that even with a limited number of images, our strategy yields a significant improvement in mesh quality. Furthermore, despite MVSNeRF [2] leveraging deep features to enable generalizable across-scene training, our method still surpasses MVSNeRF in terms of view synthesis quality, regardless of the number of images used. The mesh quality of MVSNeRF drops down since the surface estimation is biased. Note that we compare the raw mesh obtained by marching cubes without any further cleaning to the ground truth mesh.
ScanID w/o vis check w/ vis check
scan24 12.00 15.92
scan37 12.74 13.87
scan40 11.16 15.67
scan55 13.52 17.10
scan63 14.94 23.35
scan65 14.10 19.52
scan69 12.55 14.94
scan83 18.75 24.10
scan97 13.02 17.55
scan106 11.62 13.62
scan114 13.77 18.52
scan118 14.86 19.57
scan122 14.99 18.57
mean 13.69 17.87
Table 4: Ablation Study for the effectiveness of the visibility check. We compare the view synthesis results of enabling visibility and disabling visibility when reconstructing 13 scenes from DTU dataset [10] using 3 or 5 images for each scene.
Table 3 presents the performance of RayAug on the Blender dataset which outperforms NeuS by a considerable margin. Compared with NeuS under different numbers of images, our method obtains better results on all 8 scenes, demonstrating the effectiveness of our proposed method.
5.3.2 Qualitative
We also present qualitative comparisons of our method on the DTU dataset. Figure 4 shows the results for novel view synthesis, and our rendered images preserve more details compared to other methods. The rendered images using our method are more natural due to the more radiance supervision signal predicted from unseen views. As we can see from the last row, the cymbals of the drums are brighter, which is more identical to the ground truth when using our method. The results show that our method can alleviate the floaters in the reconstructed scenes. In Figure 5, our method demonstrates satisfactory performance in reconstructing the accurate depth range. The mesh reconstruction results are also provided in Figure 6.
5.3.3 Ablation Study
We investigate the effectiveness of the visibility check and its impact on the view synthesis results of ray augmentation. Table 4 shows the results of view synthesis with and without visibility check for each reconstructed scene. We train 13 scenes from the DTU dataset and set $\alpha=0.1$, which is 3 or 5 images for each scene, depending on the original number of images. The results indicate that the visibility check plays a crucial role in the ray augmentation process since augmenting rays with occluded rays would lead to wrong estimations of novel radiance, thus significantly decreasing the performance of view synthesis.
6 Conclusion
Figure 7: Unsuccessful cases in our reconstruction. It reveals the mesh at the back of the objects have several rough regions due to lack of supervisions.
In this paper, we introduce RayAug, a novel approach to address the challenge of sparse view reconstruction. Our method enhances the training of radiance fields by employing a geometry-guided strategy: checking the visibility of augmented ray directions using coarse geometry on hand. This Visibility check prevents the introduction of incorrect rays that penetrate known surfaces, leading to noisy labels, and ensures a more physically accurate augmentation process. Moreover, the radiance map incorporates a prior of natural radiance distribution, composed of a series of harmonic functions, providing confident novel radiance and proving to be more accurate than merely assigning the color from a single image. Furthermore, RayAug can be easily integrated into any surface reconstruction technique based on volume rendering. Comprehensive experiments demonstrate our method outperforms existing methods. We believe that our techniques have the potential to become a standard tool for data augmentation in implicit representation reconstruction.
However, our work also has some limitations. Firstly, while we exploit ray augmentations to tackle the sparse view problem, regions that fail the visibility check cannot be augmented, such as the back of an object (See Figure 7). Future work could consider incorporating priors by training on a larger dataset. Additionally, in the extreme case of a single view, our spherical harmonics fitting would not be useful since there is only one observation for each input pixel. Exploring alternative solutions for such cases remains an exciting avenue for future research.

### v5

@@ 4.2. Visibility Check
4.3. Radiance Map
After the visibility check, rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$ are ensured to be observed from at least one camera viewpoint. The next question is what’s the proper radiance value of rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$? In previous ray augmentation methods (Zhang et al., 2022), a surface ray has only one known radiance value, since the source view visibility check is not performed, thus they simply assign a color from one input view, which neglects the view-dependent effect and leads to inconsistent pseudo label. However, as shown in the previous discussion, a point in the scene surface could be seen from multiple views and we actually have more information regarding the radiance of surface point $\mathbf{v}$. Thus, a main contribution of our method illustrated as follows is that we utilize all reasonably credible information to reconstruct the radiance distribution of surface point $\mathbf{v}$ to infer the radiance of novel rays. Actually, not only did we fit the radiance distribution but also we achieved extrapolation using spherical harmonics expansion fitting.
@@ 4.3. Radiance Map
(4) $\mathbf{c_{v}}(\mathbf{d})=\sum_{\ell=0}^{\ell_{\mathrm{max}}}\sum_{m=-\ell}^{\ell}k_{\ell}^{m}Y_{\ell}^{m}(\mathbf{d}),$
@@
(5) $\textbf{S}_{\mathbf{v}}=\{(\mathbf{c}_{j},\mathbf{d}_{j})\}=\{(\textbf{c}_{\textbf{v},j},\mathrm{Norm}(\textbf{o}_{j}-\textbf{v}))\}_{j=1}^{N_{\mathrm{vis}}},$
@@
(6) $\mathop{\arg\min}\limits_{k_{\ell}^{m}}\>\sum_{j=1}^{N_{vis}}|\mathbf{c_{v}}(\mathbf{d}_{j})-\mathbf{c}_{j}|^{2},$
@@
4.4. Depth Warping
To best utilize the available information in the known view, we propose to propagate the depth information to other views through image warping. For pixel $p_{i}(x_{i},y_{i})$ in reference view $I_{ref}$, the corresponding pixel $p_{j}(x_{j},y_{j})$ in the $j^{th}$ unseen view $I_{unseen}$ can be formulated as
(7) $p_{j}=K_{{unseen}}T\left(K_{{ref}}^{-1}Z_{i}p_{i}\right),$
where $Z_{i}$ is the available depth of reference view, $T$ refers to the relationship between camera extrinsic matrices from $I_{{ref}}$ to $I_{{unseen}}$, and $K_{{ref}}$ and $K_{{unseen}}$ refer to the camera intrinsic matrices. We further adopt the Painter’s Algorithm when multiple points in the reference view are projected to the same point in the unseen view and select the point with the smallest depth as the warping result.
Through image warping, we can obtain a depth map of an unseen view, which can serve as a pseudo ground truth. Nevertheless, there is still an unavoidable gap between this pseudo ground truth and the real correspondence, since small misalignment in the predicted depth map can cause large errors when projected to other views. Moreover, it is quite common that the projected results contain some uncertain regions due to occlusion. To regularize the uncertain regions in the warped results, we utilize the self-supervised inverse depth smoothness loss, which uses the second-order gradients of the RGB pixel value to encourage the smoothness of the predicted depths:
(8) $\mathcal{L}_{{smooth}}\left(d_{i}\right)=e^{-\nabla^{2}\mathcal{I}\left(\mathbf{x}_{i}\right)}\left(\left|\partial_{xx}d_{i}\right|+\left|\partial_{xy}d_{i}\right|+\left|\partial_{yy}d_{i}\right|\right),$
where $d_{i}$ is the depth map, $\nabla^{2}\mathcal{I}\left(\mathbf{x}_{i}\right)$ refers to the Laplacian of pixel value at location $x_{i}$.
5. Experiments
Table 1. Quantitative comparison with selected methods on the KITTI dataset. The best and the second best results are shown in bold and underlined forms, respectively.
Dense Sparse
[…]
NeRF (Mildenhall et al., 2021) 20.03 0.652 0.502 0.203 15.07 0.552 0.603 0.253
NSG (Ost et al., 2021) 20.56 0.664 0.482 0.192 15.54 0.564 0.582 0.244
pixelNeRF (Yu et al., 2021) 19.48 0.631 0.518 0.213 14.56 0.534 0.614 0.261
SUDS (Turki et al., 2023) 20.14 0.643 0.493 0.198 15.12 0.547 0.598 0.248
MARS (Wu et al., 2023) 20.43 0.658 0.478 0.187 15.38 0.558 0.573 0.237
Urban-NeRF (Rematas et al., 2022) 20.72 0.678 0.457 0.184 15.76 0.572 0.553 0.226
MipNeRF-360 (Barron et al., 2022) 21.99 0.692 0.437 0.088 16.93 0.589 0.498 0.144
NeRF++ (Zhang et al., 2020) 20.29 0.520 0.585 3.917 17.60 0.535 0.562 4.960
Instant-NGP (Müller et al., 2022) 20.51 0.630 0.460 0.507 15.44 0.499 0.536 0.793
Ours 22.52 0.711 0.401 0.087 19.04 0.672 0.351 0.092
5.1. Novel Radiance Predicting from Spherical Harmonics
(9) $\mathbf{\hat{c}}=w\mathbf{c}_{0}+(1-w)\mathbf{c}_{1},$
(10) $w=\frac{d(\mathbf{\hat{v}},\mathbf{v}_{1})}{d(\mathbf{v}_{1},\mathbf{\hat{v}})+d(\mathbf{\hat{v}},\mathbf{v}_{2})},$
(11) $d(\mathbf{v}_{1},\mathbf{v}_{2})=\arctan\frac{\left|\mathbf{v}_{1}\times\mathbf{v}_{2}\right|}{\mathbf{v}_{1}\cdot\mathbf{v}_{2}},$
We generate a set of around $10,000$ points with all of their visible colors from the KITTI dataset (Geiger et al., 2012b) for evaluating the two methods for radiance prediction. The radiance from different views of each point is divided into two training (known) and testing (unknown) views with ratios from $8:2$ to $5:5$, in order to evaluate the performance under the different number of novel views. We report the average normalized Mean Square Error (MSE) in the novel views in Table 2. The average MSE for SH fitting is less than that of interpolation, especially when the number of novel views exceeds the number of known views. A visualization case is shown in Figure 3.
Figure 3. Visualization of radiance fitting. We project the radiance value of points distributed on the upper hemisphere into this circle to get a visualization of the fitted radiance distribution. The first row of images depicts the SH fitting results, while the second row showcases the outcomes of interpolation. It reveals that the SH fitting provides more informative radiance prediction results on the top line, while preserving the smoothness of distribution.
Table 2. MSE in the novel views between interpolation and SH fitting with different ratios between known rays and unknown rays.
@@ 5.1. Novel Radiance Predicting from Spherical Harmonics
Interpolation 0.0652 0.0744 0.0835 0.0951
SH Fitting 0.0602 0.0686 0.0779 0.0881
5.2. Implementation Details
Table 3. Evaluation and comparison of various methods on the NuScenes dataset. The best and the second best results are shown in bold and underlined forms, respectively.
Dense Sparse
@@ 5.2. Implementation Details
NeRF (Mildenhall et al., 2021) 25.32 0.798 0.505 0.190 19.88 0.762 0.535 0.212
NSG (Ost et al., 2021) 26.10 0.810 0.498 0.185 20.34 0.770 0.528 0.208
pixelNeRF (Yu et al., 2021) 26.58 0.820 0.490 0.182 20.76 0.776 0.521 0.204
SUDS (Turki et al., 2023) 27.01 0.828 0.483 0.178 21.07 0.782 0.516 0.202
MARS (Wu et al., 2023) 27.45 0.835 0.475 0.173 21.50 0.790 0.510 0.198
Urban-NeRF (Rematas et al., 2022) 27.92 0.843 0.461 0.169 20.09 0.802 0.504 0.195
MipNeRF-360 (Barron et al., 2022) 28.30 0.850 0.463 0.165 18.04 0.788 0.508 0.197
Instant-NGP (Müller et al., 2022) 27.80 0.840 0.472 0.170 16.42 0.780 0.512 0.243
Ours 29.50 0.862 0.430 0.160 26.92 0.868 0.450 0.182
Table 4. Evaluation and comparison of various methods on the Argoverse dataset. The best and the second best results are shown in bold and underlined forms, respectively.
Dense Sparse
[…]
NeRF (Mildenhall et al., 2021) 26.54 0.812 0.491 0.185 21.44 0.785 0.521 0.205
NSG (Ost et al., 2021) 27.12 0.823 0.480 0.175 22.02 0.795 0.517 0.203
pixelNeRF (Yu et al., 2021) 27.64 0.831 0.477 0.175 22.63 0.805 0.502 0.196
SUDS (Turki et al., 2023) 27.88 0.835 0.462 0.165 23.01 0.814 0.495 0.190
MARS (Wu et al., 2023) 27.92 0.843 0.469 0.162 23.50 0.826 0.485 0.185
Urban-NeRF (Rematas et al., 2022) 28.00 0.845 0.445 0.155 23.86 0.825 0.482 0.180
MipNeRF-360 (Barron et al., 2022) 29.35 0.855 0.446 0.120 25.81 0.829 0.468 0.139
Instant-NGP (Müller et al., 2022) 28.07 0.847 0.450 0.493 22.18 0.816 0.494 0.593
Ours 30.25 0.865 0.430 0.110 29.50 0.874 0.450 0.130
Dataset In our experiments, we utilized three datasets known for their extensive capture of real-world driving scenarios: KITTI (Geiger et al., 2012b), Argoverse (Chang et al., 2019), and NuScenes (Caesar et al., 2020). These datasets are challenging for 3D reconstruction due to their object movement and lighting variations. From KITTI, we selected sequences with diverse driving conditions and reduced the data to simulate a 2.5 Hz capture frequency, using 25% of available frames. This subset is henceforth referred to as the ’sparse’ dataset. In contrast, utilizing the entirety of the training set corresponds to a ’dense’ data scenario. Similarly, for Argoverse, we halved the data to match its 5 Hz frequency. NuScenes provided additional complexity with its annotated 3D bounding boxes under various conditions. For each dataset, we reserved every tenth frame for testing and used the rest for training to assess our model’s ability to handle sparse viewpoints effectively. The poses for image alignment were taken directly from the datasets’ provided odometry and tracking data to maintain consistency with the actual scale of the scenes.
@@
Implementation Details In general, we implement our ray augmentation strategy based on NeuS (Wang et al., 2021). Previous work (Müller et al., 2022) has shown that the input encoding scheme is important for reconstructing an implicit representation. Thus, we compare and choose the frequency encoding (Wang et al., 2022) for the position input, and sphere harmonics function encoding (Fridovich-Keil et al., 2022) for the direction. We find that by replacing the frequency encoding with SH function encoding for direct input, the rendered color can converge faster.
@@
Evaluation Metrics. To validate the photorealism of our synthesized views against the ground truth, we employ widely accepted metrics from the field of novel view synthesis. Specifically, we measure the Peak Signal-to-Noise Ratio (PSNR), which reflects the reconstruction accuracy in terms of image pixel intensities. The Structural Similarity Index Measure (SSIM) assesses the perceived quality of the synthesized images, accounting for texture and structural integrity. Additionally, the Learned Perceptual Image Patch Similarity (LPIPS) metric evaluates the similarity between synthesized and ground truth images based on deep features, providing an estimate of perceptual likeness.
For the evaluation of depth reconstruction quality, we follow established precedents and include the Mean Absolute Relative Error (ABSREL) and the Root Mean Squared Error (RMSE). These depth accuracy metrics provide a quantitative measure of the disparity between the estimated depth maps and the ground truth, with ABSREL focusing on the relative difference and RMSE giving the Euclidean distance error.
Figure 4. Visual comparison of scene reconstruction on the NuScenes dataset, contrasting HarmonicNeRF with MipNeRF-360 (Barron et al., 2022).
5.3. Experiment Results
Quantitative Analysis. We conduct quantitative comparisons across the KITTI (Geiger et al., 2012b), NuScenes (Caesar et al., 2020), and Argoverse (Chang et al., 2019) datasets to evaluate the performance of our method against several established baselines, including NeRF (Mildenhall et al., 2021), NSG (Ost et al., 2021), pixelNeRF (Yu et al., 2021), SUDS (Turki et al., 2023), MARS (Wu et al., 2023), Urban-NeRF (Rematas et al., 2022), MipNeRF-360 (Barron et al., 2022), and Instant-NGP (Müller et al., 2022). In the dense scenario, our method consistently outperforms all baselines, achieving the highest PSNR and SSIM scores, while maintaining the lowest LPIPS and ABSREL scores across almost all three datasets. This indicates our approach’s superior capability in synthesizing photorealistic views and reconstructing depth with high fidelity to the actual driving scenes.
The advantage of our method becomes even more pronounced in the sparse setting, where the limitations of existing techniques are more apparent due to reduced input data. Specifically, on the KITTI dataset, our method surpasses the second-best performing method, MipNeRF-360, by a significant margin in terms of PSNR (19.04 vs. 16.93) and SSIM (0.672 vs. 0.589), highlighting its effectiveness in dealing with sparse data. Similar trends are observed on the NuScenes and Argoverse datasets, where our method demonstrates exceptional performance, particularly in terms of PSNR and SSIM, further underscoring its robustness and adaptability to various driving scenarios. These results not only underscore the efficacy of HarmonicNeRF in enhancing the quality of synthesized views and the accuracy of depth reconstructions under both data-rich and data-sparse conditions but also set a new benchmark for future research in 3D scene reconstruction from sparse views.
Figure 5. Qualitative comparison of 3D scene reconstructions from the KITTI dataset using different neural radiance field methods. The top row presents original images from diverse urban settings with varying levels of detail and complexity. HarmonicNeRF, which consistently provides the clearest and most accurate depictions, with crisp textures and fine details, effectively handling challenging lighting and occlusions, and showcasing a marked improvement in both the fidelity and photorealism of the reconstructed scenes.
Qualitative Comparisons. In addition to quantitative benchmarks, qualitative assessments on the KITTI dataset showcase the visual enhancements achieved with HarmonicNeRF. The comparison in Figure 5 reveals that, while Instant-NGP and MipNeRF-360 grapple with artifacts and blurring, particularly in motion-affected areas, HarmonicNeRF produces reconstructions with remarkable clarity and detail. Our method demonstrates superior texture fidelity and depth accuracy, capturing the nuances of urban driving scene complexity with greater photorealism.
On the NuScenes dataset, the superiority of HarmonicNeRF is further evidenced through visualizations presented in Figure 4. HarmonicNeRF excels in reconstructing densely populated urban areas, preserving distinct boundaries between objects, as shown on the left side of each comparison. In contrast, the right side displays results from MipNeRF-360, which, while competent, blends complex structures together—noticeably, trees meld with vehicles, and the finer architectural features are lost to blurring.
Ablation Study. We dissect the components of HarmonicNeRF to understand their individual contributions, especially in sparse data contexts where the intricacies of each module are crucial. The ablation study results in Figure  6 highlight the impact of the key features under both dense and sparse conditions.
Without Spherical Harmonics. Replacing spherical harmonics with linear interpolation caused a noticeable performance drop. In sparse scenarios, this led to a substantial decrease in PSNR (e.g., by 2.8) and SSIM (e.g., by 0.03), validating the efficacy of spherical harmonics in handling limited data.
Without Visibility Checking. Without this component, our model was unable to effectively handle occlusions, with a significant increase in LPIPS (e.g., by 0.05) and a more pronounced decrease in SSIM in sparse data (e.g., by 0.04), confirming the crucial role of visibility checking in sparse view synthesis.
Alternative Ray Casting Methods. Substituting our surface-guided ray casting with random ray casting resulted in less accurate reconstructions, particularly under sparse conditions, where PSNR dropped by 3.0 and LPIPS increased by 0.04.
Without Depth Warping. Removing depth warping adversely affected depth estimation accuracy, especially in sparse setups, with ABSREL rising by 0.1, which points to the depth warping’s substantial role in enhancing depth precision.
Figure 6. Ablation study results on the KITTI dataset, illustrating the impact of each component under dense and sparse conditions.
6. Conclusion
In this paper, we presented HarmonicNeRF addressing the significant challenge of sparse view reconstruction in the context of autonomous driving. Our approach innovatively enhances NeRF by integrating a geometry-guided ray augmentation strategy. This strategy not only employs a visibility check to filter out non-contributory rays using coarse geometry but also leverages spherical harmonics to model the natural distribution of radiance. This dual approach ensures physically plausible augmentations and superior surface reconstruction accuracy by effectively utilizing the limited observational data typical of unbounded outdoor scenes encountered in autonomous driving.
Our comprehensive experiments, spanning the KITTI, Argoverse, and NuScenes datasets, underscore HarmonicNeRF’s ability to outperform existing methods significantly. By synthesizing novel depth views and reconstructing scenes with unprecedented precision, our method demonstrates its potential to set new standards for 3D scene reconstruction in autonomous driving applications.
However, limitations exist in our current framework. For instance, regions occluded in all observational views, such as the backside of objects, remain a challenge due to our visibility check’s inability to augment such areas. Future endeavors could explore the integration of global scene priors or learning-based methods to infer these occluded regions.